OpenStudy (anonymous):

tthe equation 3x^2+2x^2=0 represents a a. circle b. parabola c. pair of lines d.none of the above

OpenStudy (anonymous):

parabola

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

second degree polynomial in one variable.

OpenStudy (ash2326):

\[3x^2+2x^2=0 \] \[5x^2=0\] \[x=0\] x=0 is a vertical line

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=5x%5E2%3D0

OpenStudy (anonymous):

why not a pair of line...h^2-ab>0

OpenStudy (anonymous):

here, it is in one variable only(h^2) and not includes another variable(ab)

hartnn (hartnn):

is it? what are the values of a,b,and h u got?

hartnn (hartnn):

or is the question 3x^2+2y^2= 0 ??

OpenStudy (anonymous):

im sry the question i wrote is wrong it is 5y^2 not x

OpenStudy (anonymous):

then circle

OpenStudy (anonymous):

dude are u sure coz even i think its a circle